Oppenheimer Bank is offering a 30-year mortgage with an EAR of 6.375%. If you plan to borrow $160,000, what will your monthly payment be? Your monthly payment will be s(Round to the nearest cent)
Answer #1

Answer:

Given EAR =6.375%

Let r be the monthly interest rate

So (1+r)^12-1=EAR

r=(EAR+1)^(1/12)-1

r=(1+6.375%)^(1/12)-1

r=0.516%

P=$1,60,000

n=30*12=360 months

So per month payment =P*r/(1-(1+r)^-360)=160000*0.516%/(1-(1+0.516%)^-360)=$979.12